MR=6+10x-18x^2



X=1



Sh 4000



Find total revenue

1
Expert's answer
2022-01-13T15:54:26-0500

"MR=6+10x-18x\u00b2"

("MR" represents Marginal Revenue)

"MR=\frac{dR}{dx}"

("R" represents Total revenue)

Integrating both sides,we have

"\int{MR} dx+k=R"

"\implies R=\int{MR}dx+ k"

"=\int(6+10x-18x²)dx+k"

Using sum rule,

"=\int6dx+\int10xdx-\int18x²dx+k"

"=6x+5x\u00b2-6x\u00b3+k"

Since "R=0" when "x=0" ,then "k=0"

"\implies R=6x+5x²-6x³"